OpenStudy (anonymous):

A large totem pole near Kalama, Washington, is 121 feet tall. On a particular day at noon it casts a 213 foot shadow. What is the sun's angle of elevation at the time ? Round to the nearest tenth

OpenStudy (kropot72):

 Let the angle of elevation be theta. Then we have: \[\tan \theta=\frac{121}{213}\] from which we find theta as: \[\theta=\tan^{-1} (\frac{121}{213})\]

OpenStudy (anonymous):

wait so what would be the answer ?

OpenStudy (kropot72):

Just use your calculator to get the answer. If you haven't got a calculator, then use one of the online calculators.

OpenStudy (kropot72):

You can find a suitable calculator here: http://www.rapidtables.com/calc/math/Arctan_Calculator.htm

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)